# DRUPI: Dataset Reduction Using Privileged Information

### v2

@@ 5 Discussion
Different ways for synthesizing privileged information. As discussed in Section 3.2, there are multiple methods for synthesizing feature labels. A straightforward approach is to assign feature labels using a pre-trained model. However, as shown in Figure 4, this can sometimes degrade dataset quality. In contrast, learning feature labels offers greater flexibility and adaptability. As demonstrated in Figure 4(a), reduced datasets with learned feature labels significantly outperform those with directly assigned features. This is because feature labels extracted from a pre-trained model often lead to overly discriminative features with low diversity. The empirical results also support the observation that overly discriminative feature labels with strong task supervision can hurt performance in Figure 3. More detailed results are provided in Table 10 in Appendix B.3.
@@
Impact of feature label versatility and methods for utilizing feature labels. We investigated the effect of synthesizing multiple feature labels for a single data-label pair. As shown in Figure 4(b), experiments demonstrate that increasing the number of feature labels enhances performance, likely due to the greater versatility captured by additional features. However, too many feature labels for a single input can introduce excessive diversity, leading to degraded performance. This verifies the trade-off between the diversity and discriminability of feature labels. Furthermore, averaging multiple feature labels outperforms random selection, which enables us to save only the averaged feature labels. Hence, increasing the number of feature labels does not bring more storage overhead.
@@
For more details on different methods of initializing feature labels, please refer to Appendix B.4. Ablation studies on the regression magnitude of MSE are provided in Appendix C.1. Additionally, we explored the potential of feature regression using losses beyond MSE in Appendix C.2.
@@ 6 Conclusion
In this paper, we introduced DRUPI, a novel framework that synthesizes privileged information for reduced datasets. To the best of our knowledge, DRUPI is the first approach to go beyond the traditional data-label paradigm by utilizing synthesized feature labels. Extensive experiments on ImageNet, CIFAR-10/100, and Tiny ImageNet validate the effectiveness of DRUPI, demonstrating significant improvements in model performance when integrated with existing reduction techniques. Additionally, we showed that achieving a balance between the discriminability and diversity of the synthesized feature labels is crucial for maximizing the quality of the reduced dataset.
@@ A.1 Computational Resources
The training was conducted on NVIDIA GPUs, specifically RTX 4090 and A100. All coreset selection experiments were run on A100 GPUs, while all DC-related experiments were carried out with RTX 4090. For MTT and DATM experiments on CIFAR-100 with IPC 10 and 50, as well as Tiny ImageNet, we utilized four NVIDIA A100 GPUs. Results of smaller datasets and lower IPC settings were conducted with RTX 4090.
@@ A.2 Coreset Selection
We first employed several coreset selection methods to initialize our reduced dataset, specifically using Herding (Welling, 2009), k-Center (Har-Peled & Mazumdar, 2004), and Forgetting (Toneva et al., 2018), where each data point was selected based on scores from a pre-trained ResNet-18. Next, we synthesized feature labels for the coreset using DC Zhao et al. (2020), assigning these labels to the intermediate features of a ConvNet trained for just one epoch. We also fine-tuned the initial images using the same method as for feature label learning, although we found that simply synthesizing feature labels could already bring performance improvement for reduced datasets.
@@ A.3 Dataset Distillation
For dataset distillation methods, we initialized the reduced datasets (images and labels) using distilled datasets and applied the same distillation method to synthesize feature labels. Specifically, we used DC (Zhao et al., 2020), MTT (Cazenavette et al., 2022), and DATM (Guo et al., 2023) for both data-label initialization and feature label synthesis.
We followed the original image-label synthesis settings of these distillation methods to generate feature labels. Table 6 provide the detailed hyperparameter settings used for experiments on CIFAR-10, CIFAR-100, Tiny ImageNet. Table 7 provides the parameter settings for the MTT method on the ImageNet subsets. By default, we set $n_{feat}$ to 1 and $\lambda_{reg}$ to 0.01. We explored different configurations of Images Per Class (IPC), specifically IPC = {1, 10, 50} for CIFAR-10 and CIFAR-100, and IPC = {1, 10} for Tiny ImageNet.
@@
Dataset IPC Method $\lambda_{reg}$ $\lambda_{task}$ $n_{feat}$
CIFAR-10 1 DC 1.5 0.1 1
@@
Dataset IPC $\lambda_{reg}$
ImageNette 1 0.005
@@
For each method, the hyperparameters for feature synthesis are fine-tuned across different datasets and IPC settings to achieve optimal performance. For instance, in CIFAR-10 (1 IPC), the DC method utilizes $\lambda_{reg}=1.5$, $\lambda_{task}=0.1$, and $n_{feat}=1$. In contrast, for 50 IPC, the same method adjusts its hyperparameters to $\lambda_{reg}=0.001$, $\lambda_{task}=0.005$, and $n_{feat}=1$. Similar fine-tuning is performed for all datasets and IPC values across each method.
@@
Appendix B Additional Results on DRUPI
@@ B.1 Further Performance Results
In Section 4.2, we presented results for two dataset distillation methods, DC and MTT. Here, we provide the results for DATM on the CIFAR-10 and CIFAR-100 datasets. We initialized the images and labels with DATM and used it to synthesize feature labels for the reduced dataset. Experiments were conducted using a ConvNet for both distillation and evaluation tasks. We additionally provide baselines such as (Zhou et al., 2022; Loo et al., 2023; Cui et al., 2023; Du et al., 2023; Guo et al., 2023) .Table 8 shows the performance improvements of DRUPI over various existing methods under different IPC settings. Specifically, the datasets used in these experiments include CIFAR-10 with IPC = {1, 10, 50} and CIFAR-100 with IPC = {1, 10}. Each dataset was further evaluated with varying data fractions, allowing us to assess the generalization of the methods across different data availability scenarios.
Table 8: The application of DRUPI to DATM across CIFAR-10, CIFAR-100, accompanied by a comparative analysis with existing methods. ConvNet is utilized for both distillation and evaluation. Our methodology demonstrates enhanced performance compared to previous results.The ↑ symbol signifies performance enhancements compared to random selection.
@@
Random 15.4$\pm$0.3 31.0$\pm$0.5 50.6$\pm$0.3 4.2$\pm$0.3 14.6$\pm$0.5
KIP 49.9$\pm$0.2 62.7$\pm$0.3 68.6$\pm$0.3 15.7$\pm$0.2 28.3$\pm$0.1
FRePo 46.8$\pm$0.7 65.5$\pm$0.4 71.7$\pm$0.4 28.7$\pm$0.2 42.5$\pm$0.4
RCIG 53.9$\pm$1.0 69.1$\pm$0.4 73.5$\pm$0.3 39.3$\pm$0.4 44.1$\pm$0.4
DM 26.0$\pm$0.8 48.9$\pm$0.6 63.0$\pm$0.4 11.4$\pm$0.3 29.7$\pm$0.3
DSA 28.8$\pm$0.7 52.1$\pm$0.5 60.6$\pm$0.5 13.9$\pm$0.3 32.3$\pm$0.3
DCC 32.9$\pm$0.8 49.4$\pm$0.5 61.6$\pm$0.4 13.3$\pm$0.3 30.6$\pm$0.4
DSAC 34.0$\pm$0.7 54.5$\pm$0.5 64.2$\pm$0.4 14.6$\pm$0.3 33.5$\pm$0.3
CAFE 30.3$\pm$1.1 46.3$\pm$0.6 55.5$\pm$0.6 12.9$\pm$0.3 27.8$\pm$0.3
IDM 45.6$\pm$0.7 58.6$\pm$0.1 67.5$\pm$0.1 20.1$\pm$0.3 45.1$\pm$0.1
TESLA 48.5$\pm$0.8 66.4$\pm$0.8 72.6$\pm$0.7 24.8$\pm$0.4 41.7$\pm$0.3
FTD 46.0$\pm$0.4 65.3$\pm$0.4 73.2$\pm$0.2 24.4$\pm$0.4 42.5$\pm$0.2
DATM 46.9$\pm$0.5 66.8$\pm$0.2 76.1$\pm$0.3 27.9$\pm$0.2 47.2$\pm$0.4
+DRUPI 48.0$\pm$0.4 67.8$\pm$0.3 76.4$\pm$0.3 28.4$\pm$0.5 47.6$\pm$0.2
@@
Full Dataset 84.8$\pm$0.1 56.2$\pm$0.3
B.2 Results on Cross-Architecture Evaluation with Attention Labels
In addition to the results presented in Table 4 and 5 of the main paper, which summarize the performance of utilizing the synthesized datasets from pruning and dataset distillation to initialize the feature labels, we provide additional results in Table 9 to evaluate the cross-architecture generalization by using both feature labels and attention labels on different reduced datasets.
Table 9 illustrates the cross-architecture evaluations conducted for both dataset condensation (DC) and MTT across various network architectures. In these experiments, we propose DRUPI-F (feature labels learned by DC) and its variant DRUPI-A, which is further enhanced by pooling through spatial attention. Similarly, feature labels synthesized with MTT are also average pooled into channel attention labels. We utilized several networks for cross evaluatoin, including LeNet, MLP, ResNet, VGG, ConvNet, and AlexNet.
The table demonstrates that initializing the reduced dataset with feature labels assigned through intermediate features of pre-trained networks leads to significant performance improvements across all architectures. Further gains are observed when channel attention pooling is applied. By leveraging Eq. (5) to learn and update the feature labels, the reduced dataset consistently yields competitive or improved performance across different networks. This highlights the benefits of feature and attention labels in enhancing model generalization, with particularly notable improvements observed in ConvNet and AlexNet settings. Channel attention pooling further contributes to these gains, reinforcing the effectiveness of this approach across architectures.
LeNet MLP ResNet VGG ConvNet AlexNet
DC 16.8$\pm$5.2 28.0$\pm$0.6 36.2$\pm$1.6 35.4$\pm$0.6 44.5$\pm$0.5 20.1$\pm$4.5
+DRUPI-F 24.7$\pm$4.3 32.3$\pm$1.0 37.5$\pm$1.8 35.8$\pm$0.8 47.1$\pm$0.9 27.3$\pm$3.2
+DRUPI-A 23.7$\pm$6.6 27.9$\pm$0.5 37.7$\pm$1.1 35.6$\pm$0.7 45.6$\pm$0.5 25.3$\pm$3.0
MTT 29.1$\pm$1.5 29.1$\pm$0.4 35.1$\pm$1.1 31.4$\pm$0.9 45.6$\pm$0.7 24.0$\pm$0.8
+DRUPI-A 31.6$\pm$1.7 29.4$\pm$0.5 36.7$\pm$5.2 37.9$\pm$2.8 47.0$\pm$0.7 34.1$\pm$4.4
B.3 Additional Results on Different Methods for Synthesizing Privileged Information
As a complement to Figure 4(a), Table 10 provides a more comprehensive overview of the experimental results on different methods for synthesizing privileged information. Specifically, it presents a detailed comparison between direct feature label assignment and the DRUPI framework across three dataset distillation methods: DC, MTT, and DATM. These evaluations were conducted on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets with varying images per class (IPC) values, and include results for different fractions of the full dataset.
For each method, the results are provided for various IPC settings, showing the performance both with direct feature assignment and with the inclusion of DRUPI (feature label). The DRUPI framework consistently improves performance across different methods and datasets, as indicated by the higher accuracy values. Specifically, DRUPI shows substantial improvements over direct assignment, particularly in the low-data regime, such as CIFAR-100 1 IPC and Tiny ImageNet 1 IPC. These findings underscore the effectiveness of DRUPI in enhancing the generalization and performance of dataset distillation techniques, even with limited data.
Table 10: Experimental results comparing direct feature label assignment and the DRUPI framework across three dataset distillation methods (DC, MTT, DATM) on various datasets with different IPC values.
@@ B.3 Additional Results on Different Methods for Synthesizing Privileged Information
DC 28.3$\pm$0.5 44.9$\pm$0.5 53.9$\pm$0.5 12.8$\pm$0.3 25.2$\pm$0.3 29.8$\pm$0.3 -
Directly Assign 28.3$\pm$0.8 45.1$\pm$0.5 54.1$\pm$0.4 12.7$\pm$0.4 25.1$\pm$0.4 29.8$\pm$0.4 -
DRUPI 31.5$\pm$0.9 47.4$\pm$0.9 55.0$\pm$0.5 14.9$\pm$0.4 29.2$\pm$0.5 30.9$\pm$0.5 -
MTT 46.2$\pm$0.8 65.4$\pm$0.7 71.6$\pm$0.2 24.3$\pm$0.3 39.7$\pm$0.4 47.7$\pm$0.2 8.8$\pm$0.3
Directly Assign 40.8$\pm$1.8 56.2$\pm$1.1 66.1$\pm$0.5 23.9$\pm$0.5 38.1$\pm$0.4 47.4$\pm$0.2 8.1$\pm$0.4
DRUPI 47.4$\pm$0.5 65.8$\pm$0.6 71.7$\pm$0.2 25.6$\pm$0.4 40.8$\pm$0.3 47.9$\pm$0.3 11.0$\pm$0.1
DATM 46.9$\pm$0.5 66.8$\pm$0.2 76.1$\pm$0.3 27.9$\pm$0.2 47.2$\pm$0.4 55.0$\pm$0.2 17.1$\pm$0.3
Directly Assign 46.3$\pm$0.7 64.5$\pm$0.5 73.7$\pm$0.4 26.6$\pm$0.3 36.2$\pm$0.5 55.5$\pm$0.3 15.6$\pm$0.1
DRUPI 48.0$\pm$0.4 67.8$\pm$0.3 76.4$\pm$0.3 28.4$\pm$0.5 47.6$\pm$0.2 55.0$\pm$0.1 17.6$\pm$0.1
Full Dataset 84.8$\pm$0.1 56.2$\pm$0.3 37.6$\pm$0.4
B.4 Effects on Feature Label Initialization
@@
Figure 5 presents a performance comparison between two feature initialization approaches—random noise and assigned features—across the CIFAR-10 and CIFAR-100 datasets with varying images per class (IPC) settings. In this comparison, noise (yellow bars) refers to randomly initialized feature labels, while real (blue bars) refers to initialization using features extracted from synthetic datasets passed through the network.
For CIFAR-10 (1 IPC) setting, the initialization with assigned features method achieves significantly higher performance (around 33%) compared to the noise-based initialization (approximately 30%). A similar trend is observed in the CIFAR-10 (10 IPC) case, where initialization with assigned features substantially outperforms noise.
On the CIFAR-100 (1 IPC) dataset, initialization with assigned features also demonstrates better performance, with a clear margin over noise initialization, where using assigned features reaches around 29% while noise remains below 28%. The same pattern holds for CIFAR-100 (10 IPC), where initialization with assigned features achieves a accruacy of over 47%, far exceeding the noise initialization’s performance of approximately 45%.
Figure 5: Comparison of noise initialization (yellow) and initialization with assigned features (blue) from a pre-trained ConvNet on CIFAR-10 and CIFAR-100 across different IPC settings.
@@ C.1 Sensitivity of Regression Magnitude
We further investigated the influence of $\lambda_{reg}$ on the model’s performance. The results in Table 11 demonstrate that varying the MSE regularization parameter $\lambda_{reg}$ does not significantly impact the final accuracy. While there is a slight improvement as $\lambda_{reg}$ increases, from 28.91% at $\lambda_{reg}=0.05$ to 30.74% at $\lambda_{reg}=10$, the overall effect remains relatively small. This suggests that the model’s performance is not highly sensitive to the regularization weight for MSE loss within the tested range, indicating that other factors may have a more dominant influence on accuracy. In this case, regularization helps prevent overfitting but does not drastically change the model’s ability to generalize within the 1 IPC setting for CIFAR-10.
Table 11: Results on CIFAR-10 1 IPC for different $\lambda_{reg}$.

### v3

@@ 5 Discussion
Different ways for synthesizing privileged information. As discussed in Section 3.2, there are multiple methods for synthesizing feature labels. A straightforward approach is to assign feature labels using a pre-trained model. However, as shown in Figure 4, this can sometimes degrade dataset quality. In contrast, learning feature labels offers greater flexibility and adaptability. As demonstrated in Figure 4(a), reduced datasets with learned feature labels significantly outperform those with directly assigned features. This is because feature labels extracted from a pre-trained model often lead to overly discriminative features with low diversity. The empirical results also support the observation that overly discriminative feature labels with strong task supervision can hurt performance in Figure 3. More detailed results about the experiments on different methods are provided in Table 11 in Appendix B.2.
@@
Impact of feature label versatility and methods for utilizing feature labels. We investigated the effect of synthesizing multiple feature labels for a single data-label pair. As shown in Figure 4(b), experiments demonstrate that increasing the number of feature labels enhances performance, likely due to the greater versatility captured by additional features. However, too many feature labels for a single input can introduce excessive diversity, leading to degraded performance. This verifies the trade-off between the diversity and discriminability of feature labels. Furthermore, averaging multiple feature labels outperforms random selection, which enables us to save only the averaged feature labels. Thus, increasing the number of feature labels introduces no extra storage overhead.
@@
Adaptive synthesis guidance applied in RDED. When applied to RDED, DCPI operates as a flexible process guided by IPC. It leverages early-epoch, high-entropy soft labelsQin et al. (2024) for scenarios with small IPC and late-epoch, full soft labels for scenarios with large IPC. For small IPC values, DCPI steers the data synthesis to either maximize diversity based on the early-stage teacher’s features (consistent with the high-entropy nature of its soft labels) or to amplify emerging discriminative signals, thereby enriching the limited data. In contrast, for large IPC settings, DCPI employs precise soft labels from fully converged teachers and prioritizes ensuring that the synthetic data demonstrates maximal discriminability within the converged teacher’s feature space, effectively converting label precision into highly informative and separable training examples.
For details on feature label initialization, see Appendix B.3. Ablation studies on MSE regression magnitude are in Table 12 and Appendix C.1. Feature regression with other losses is explored in Table 13 and Appendix C.2.
@@ 6 Conclusion
In this paper, we introduced DCPI, a novel framework that synthesizes privileged information for reduced datasets. To the best of our knowledge, DCPI is the first approach to go beyond the traditional data-label paradigm by utilizing synthesized feature labels. Extensive experiments on ImageNet, CIFAR-10/100 and Tiny ImageNet validate the effectiveness of DCPI, demonstrating significant improvements in model performance when integrated with existing condensation techniques. Additionally, we showed that achieving a balance between the discriminability and diversity of the synthesized feature labels is crucial for maximizing the quality of the reduced dataset.
@@ A.1 Computational Resources
The training was conducted on NVIDIA GPUs, specifically RTX 4090 and A100. All coreset selection experiments were run on A100 GPUs, while all DC-related experiments were carried out with RTX 4090. For MTT and DATM experiments on CIFAR-100 with fraction = 0.2% and 2%, as well as Tiny ImageNet, we utilized four NVIDIA A100 GPUs. Results of smaller datasets and lower fraction settings were conducted with RTX 4090.
@@ A.2 Coreset Selection
We first employed several coreset selection methods to initialize our reduced dataset, specifically using Herding Welling (2009), k-Center Har-Peled and Mazumdar (2004), and Forgetting Toneva et al. (2018), where each data point was selected based on scores from a pre-trained ResNet-18. Next, we synthesized feature labels for the coreset using DC Zhao et al. (2020), assigning these labels to the intermediate features of a ConvNet trained for just one epoch. We also fine-tuned the initial images using the same method as for feature label learning, although we found that simply synthesizing feature labels could already bring performance improvement for reduced datasets.
@@ A.3 Dataset Distillation
For dataset distillation methods, we initialized the reduced datasets (images and labels) using distilled datasets and applied the same distillation method to synthesize feature labels. Specifically, we used DC Zhao et al. (2020), MTT Cazenavette et al. (2022), and DATM Guo et al. (2023) for both data-label initialization and feature label synthesis.
We followed the original image-label synthesis settings of these distillation methods to generate feature labels. Table 8 provide the detailed hyperparameter settings used for experiments on CIFAR-10, CIFAR-100, Tiny ImageNet. Table 9 provides the parameter settings for the MTT method on the ImageNet subsets. By default, we set $n_{feat}$ to 1 and $\lambda_{reg}$ to 0.01. We explored different configurations of fraction, specifically fraction(%) = {0.02, 0.2, 1} for CIFAR-10, fraction = {0.2, 2, 10} for CIFAR-100 and fraction = {0.2, 2} for Tiny ImageNet.
@@
Dataset fraction(%) Method $\lambda_{reg}$ $\lambda_{task}$ $n_{feat}$
CIFAR-10 0.02 DC 1.5 0.1 1
@@
Dataset fraction $\lambda_{reg}$
ImageNette 0.1 0.005
@@
For each method, the hyperparameters for feature synthesis are fine-tuned across different datasets and fraction settings to achieve optimal performance. For instance, in CIFAR-10 (fraction = 0.02%), the DC method utilizes $\lambda_{reg}=1.5$, $\lambda_{task}=0.1$, and $n_{feat}=1$. In contrast, for fraction = 1%, the same method adjusts its hyperparameters to $\lambda_{reg}=0.01$, $\lambda_{task}=0.01$, and $n_{feat}=1$. Similar fine-tuning is performed for all datasets and fraction values across each method.
@@
Appendix B Additional Results on DCPI
@@ B.1 Further Performance Results
In Section 4.2, we presented results for two dataset distillation methods, DC and MTT. Here, we provide the results for DATM on the CIFAR-10 and CIFAR-100 datasets. We initialized the images and labels with DATM and used it to synthesize feature labels for the reduced dataset. Experiments were conducted using a ConvNet for both distillation and evaluation tasks. We additionally provide baselines such as Zhou et al. (2022); Loo et al. (2023); Cui et al. (2023); Du et al. (2023); Guo et al. (2023) .Table 10 shows the performance improvements of DCPI over various existing methods under different fraction settings. Specifically, the datasets used in these experiments include CIFAR-10 with fraction(%) = {0.02, 0.2, 1} and CIFAR-100 with fraction(%) = {0.2, 2}. Each dataset was further evaluated with varying data fractions, allowing us to assess the generalization of the methods across different data availability scenarios.
Table 10: The application of DCPI to DATM across CIFAR-10, CIFAR-100, accompanied by a comparative analysis with existing methods. ConvNet is utilized for both distillation and evaluation. Our methodology demonstrates enhanced performance compared to previous results.The ↑ symbol signifies performance enhancements compared to random selection.
@@
[…]
@@
[…]
B.2 Additional Results on Different Methods for Synthesizing Privileged Information
As a complement to Figure 4(a), Table 11 provides a more comprehensive overview of the experimental results on different methods for synthesizing privileged information. Specifically, it presents a detailed comparison between direct feature label assignment and the DCPI framework across three dataset distillation methods: DC, MTT, and DATM. These evaluations were conducted on CIFAR-10, CIFAR-100, and Tiny ImageNet datasets with varying fraction values, and include results for different fractions of the full dataset.
For each method, the results are provided for various fraction settings, showing the performance both with direct feature assignment and with the inclusion of DCPI (feature label). The DCPI framework consistently improves performance across different methods and datasets, as indicated by the higher accuracy values. Specifically, DCPI shows substantial improvements over direct assignment, particularly in the low-data regime, such as CIFAR-100(0.2%) and Tiny ImageNet(0.2%). These findings underscore the effectiveness of DCPI in enhancing the generalization and performance of dataset distillation techniques, even with limited data.
Table 11: Experimental results comparing direct feature label assignment and the DCPI framework across three dataset distillation methods (DC, MTT, DATM) on various datasets with different fraction values.
@@ B.2 Additional Results on Different Methods for Synthesizing Privileged Information
[…]
B.3 Effects on Feature Label Initialization
@@
Figure 5 presents a performance comparison between two feature initialization approaches—random noise and assigned features—across the CIFAR-10 and CIFAR-100 datasets with varying fraction settings. In this comparison, noise (yellow bars) refers to randomly initialized feature labels, while real (blue bars) refers to initialization using features extracted from synthetic datasets passed through the network.
For CIFAR-10 (IPC = 1) setting, the initialization with assigned features method achieves significantly higher performance (around 33%) compared to the noise-based initialization (approximately 30%). A similar trend is observed in the CIFAR-10 (IPC = 10) case, where initialization with assigned features substantially outperforms noise.
On the CIFAR-100 (IPC = 1) dataset, initialization with assigned features also demonstrates better performance, with a clear margin over noise initialization, where using assigned features reaches around 29% while noise remains below 28%. The same pattern holds for CIFAR-100 (IPC = 10), where initialization with assigned features achieves a accruacy of over 47%, far exceeding the noise initialization’s performance of approximately 45%.
Figure 5: Comparison of noise initialization (yellow) and initialization with assigned features (blue) from a pre-trained ConvNet on CIFAR-10 and CIFAR-100 across different fraction settings.
@@ C.1 Sensitivity of Regression Magnitude
We further investigated the influence of $\lambda_{reg}$ on the model’s performance. The results in Table 12 demonstrate that varying the MSE regularization parameter $\lambda_{reg}$ does not significantly impact the final accuracy. While there is a slight improvement as $\lambda_{reg}$ increases, from 28.91% at $\lambda_{reg}=0.05$ to 30.74% at $\lambda_{reg}=10$, the overall effect remains relatively small. This suggests that the model’s performance is not highly sensitive to the regularization weight for MSE loss within the tested range, indicating that other factors may have a more dominant influence on accuracy. In this case, regularization helps prevent overfitting but does not drastically change the model’s ability to generalize within the fraction = 0.02% setting for CIFAR-10.
Table 12: Results on CIFAR-10 (fraction = 0.02%) for different $\lambda_{reg}$.